# Beyond Classification Accuracy: Neural-MedBench and the Need for Deeper Reasoning Benchmarks

### v2

@@ 1 Introduction
Recent advances in vision-language models (VLMs) have led to striking improvements across a wide range of medical AI tasks. On standard benchmarks such as MedMNIST v2 Yang et al. (2023) and MultiMedQA Zhou et al. (2024), state-of-the-art models achieve near-human or even superhuman performance in label prediction and image–text alignment. These results have created an impression that medical VLMs are nearing clinical readiness. Yet, as the clinical reasoning literature has recently underscored Schwartzstein (2024), safe and effective diagnostic practice, especially in high-stakes fields such as neurology, demands more than classification accuracy: it requires multimodal synthesis, ambiguity resolution, and the capacity to justify conclusions in a manner consistent with clinical logic. Current benchmarks, despite their scale, rarely capture these aspects. We argue that this discrepancy creates an evaluation illusion, a misleading sense of model capability, where strong performance on shallow benchmarks obscures deep weaknesses in reasoning fidelity and clinical trustworthiness.
To address this gap, we propose a new perspective on evaluation: a Two-Axis Evaluation Framework for medical AI. Along the first axis, Breadth, large-scale datasets assess statistical generalization and coverage across populations and conditions. This is where almost all existing benchmarks operate. However, recent work such as DiagnosisArena Zhu et al. (2025) has shown that relying solely on such breadth-oriented evaluations leads to overestimation of model capability, as they fail to capture true clinical reasoning challenges. Along the second axis, Depth, compact but diagnostically complex, expert-curated benchmarks assess reasoning fidelity, forcing models to integrate multimodal signals, reason under uncertainty, and provide structured justifications, an approach that echoes trends in MedAgentsBench Tang et al. (2025). Importantly, we hypothesize that these axes are largely uncorrelated: success on breadth benchmarks does not guarantee competence in depth scenarios. This hypothesis is supported by findings that even expert LLMs struggle with self-awareness and metacognitive reasoning on challenging cases, as shown in MetaMedQA Griot et al. (2025). A complete picture of model readiness therefore requires evaluation along both dimensions.
@@
To operationalize the Depth axis in clinical neurology, we introduce Neural-MedBench, a diagnostic benchmark deliberately designed as a “stress test.” Neural-MedBench comprises 120 expert-annotated multimodal cases, including multi-sequence MRI scans, structured electronic health records (EHRs), and clinical narratives, yielding 200 reasoning-intensive tasks. Its design mirrors practices in medical education, such as Objective Structured Clinical Examinations (OSCEs), where carefully constructed cases are used to assess reasoning skills under uncertainty Geathers et al. (2025); Siegelman et al. (2024). Unlike prior resources focused on breadth, Neural-MedBench is compact by design, prioritizing reasoning density over volume, enabling cost-effective yet high-signal evaluation of clinical reasoning.
@@ 2.1 Medical VLM Benchmarks
A wide range of benchmarks have been developed for evaluating medical VLMs. Early datasets such as ChestX-ray14 Wang et al. (2017), CheXpert Irvin et al. (2019), and MIMIC-CXR Johnson et al. (2019) primarily enabled large-scale classification and report generation. Organ- or disease-specific challenges—including LUNA for pulmonary nodules Setio et al. (2017), LiTS for liver tumors Bilic et al. (2019), BraTS for brain tumors Menze et al. (2015), and ISLES for stroke Winzeck et al. (2017)—expanded evaluation to segmentation and lesion characterization. More recent multimodal benchmarks target richer tasks: ROCO v2 Rückert et al. (2024) for radiology captioning, RadGraph Jain et al. (2021) for entity–relation extraction, and MedFMC Zhang et al. (2023) for federated multi-center classification. VQA-style benchmarks, such as VQA-RAD Lau et al. (2018), PathVQA He et al. (2020), and OmniMedVQA Hu et al. (2024), test text–image reasoning but often rely on templated questions, limiting their coverage of authentic diagnostic challenges. Recent meta-benchmarks (e.g., Rad-Bench Kuo et al. (2024), DiagnosisArena Zhu et al. (2025), MedAgentsBench Tang et al. (2025)) have emphasized comprehensiveness, yet remain breadth-oriented: they focus on scale and diversity rather than probing reasoning under uncertainty.
@@ 2.2 VLM Evaluation Methodologies
Medical VLMs are typically evaluated via three strategies: reference-based, reference-free, or hybrid metrics. Traditional reference-based methods include BLEU, ROUGE, and CIDEr for surface overlap, as well as semantic similarity scores such as BERTScore Zhang et al. (2019), CLIPScore Hessel et al. (2021), and domain-specific adaptations like CheXbert Smit et al. (2020) or RadGraph-F1 Jain et al. (2021). While scalable, these methods cannot assure that generated outputs are clinically grounded. Reference-free approaches, such as saliency attribution (e.g., Grad-CAM Selvaraju et al. (2020)) or image–text alignment tests, serve as indirect proxies for reasoning but lack explanatory power. More recent frameworks have introduced LLM-based grading systems for evaluating factual accuracy and reasoning plausibility in free-form clinical responses. For example, the Expert-of-Experts Verification and Alignment (EVAL) framework streamlines clinician-led evaluation in bleeding scenarios Smith et al. (2025), while LLMEval-Med combines expert-designed checklists with LLM-as-Judge dynamics to improve reliability in medical QA and reasoning tasks Zhang et al. (2025); Zhi et al. (2025). Despite their automation advantages, these approaches remain subject to ongoing reliability validation.
@@ 4.1 Models Benchmarked
We benchmark three categories of models to provide a comprehensive view of current VLM capabilities. First, we include proprietary frontier systems with strong multimodal reasoning ability, namely GPT-4o OpenAI (2024) and Gemini Comanici et al. (2025), which serve as reference baselines. Second, we evaluate powerful open-source generalist models such as Claude Anthropic (2024), Doubao Team (2024), Qwen-VL-Plus Bai et al. (2023). Finally, to assess the effect of domain-specific adaptation, we benchmark medical-specialized VLMs including LLaVA-Med Li et al. (2023), RadFM Wu et al. (2023), Med-Flamingo Moor et al. (2023) and MedGemma Sellergren et al. (2025), which represent current efforts to tailor VLMs for clinical applications. A full list of models, versions, and hyperparameters is provided in Appendix C.
@@ 7 Ethics Details
  • Radiopaedia: Radiopaedia data used in this study are provided under the Creative Commons Attribution-NonCommercial-ShareAlike 3.0 ($C$-BY-NC-SA 3.0) license. The data were utilized for non-commercial purposes with proper attribution to Radiopaedia and are redistributed under the same licensing terms.

### v3

@@ 1 Introduction
Recent advances in vision-language models (VLMs) have led to striking improvements across a wide range of medical AI tasks. On standard benchmarks such as MedMNIST v2 yang2023medmnist and MultiMedQA zhou2024multifaceteval, state-of-the-art models achieve near-human or even superhuman performance in label prediction and image–text alignment. These results have created an impression that medical VLMs are nearing clinical readiness. Yet, as the clinical reasoning literature has recently underscored schwartzstein2024clinical, safe and effective diagnostic practice, especially in high-stakes fields such as neurology, demands more than classification accuracy: it requires multimodal synthesis, ambiguity resolution, and the capacity to justify conclusions in a manner consistent with clinical logic. Current benchmarks, despite their scale, rarely capture these aspects. We argue that this discrepancy creates an evaluation illusion, a misleading sense of model capability, where strong performance on shallow benchmarks obscures deep weaknesses in reasoning fidelity and clinical trustworthiness.
To address this gap, we propose a new perspective on evaluation: a Two-Axis Evaluation Framework for medical AI. Along the first axis, Breadth, large-scale datasets assess statistical generalization and coverage across populations and conditions. This is where almost all existing benchmarks operate. However, recent work such as DiagnosisArena zhu2025diagnosisarena has shown that relying solely on such breadth-oriented evaluations leads to overestimation of model capability, as they fail to capture true clinical reasoning challenges. Along the second axis, Depth, compact but diagnostically complex, expert-curated benchmarks assess reasoning fidelity, forcing models to integrate multimodal signals, reason under uncertainty, and provide structured justifications, an approach that echoes trends in MedAgentsBench tang2025medagentsbench. Importantly, we hypothesize that these axes are largely uncorrelated: success on breadth benchmarks does not guarantee competence in depth scenarios. This hypothesis is supported by findings that even expert LLMs struggle with self-awareness and metacognitive reasoning on challenging cases, as shown in MetaMedQA griot2025large. A complete picture of model readiness therefore requires evaluation along both dimensions.
@@
To operationalize the Depth axis in clinical neurology, we introduce Neural-MedBench, a diagnostic benchmark deliberately designed as a “stress test.” Neural-MedBench comprises 120 expert-annotated multimodal cases, including multi-sequence MRI scans, structured electronic health records (EHRs), and clinical narratives, yielding 200 reasoning-intensive tasks. Its design mirrors practices in medical education, such as Objective Structured Clinical Examinations (OSCEs), where carefully constructed cases are used to assess reasoning skills under uncertainty geathers2025benchmarking; siegelman2024assessment. Unlike prior resources focused on breadth, Neural-MedBench is compact by design, prioritizing reasoning density over volume, enabling cost-effective yet high-signal evaluation of clinical reasoning.
@@ 2.1 Medical VLM Benchmarks
A wide range of benchmarks have been developed for evaluating medical VLMs. Early datasets such as ChestX-ray14 wang2017chestx, CheXpert irvin2019chexpert, and MIMIC-CXR johnson2019mimic primarily enabled large-scale classification and report generation. Organ- or disease-specific challenges—including LUNA for pulmonary nodules setio2017validation, LiTS for liver tumors bilic2019liver, BraTS for brain tumors menze2015multimodal, and ISLES for stroke winzeck2018isles—expanded evaluation to segmentation and lesion characterization. More recent multimodal benchmarks target richer tasks: ROCO v2 ruckert2024rocov2 for radiology captioning, RadGraph jain2021radgraph for entity–relation extraction, and MedFMC zhang2023medfmc for federated multi-center classification. VQA-style benchmarks, such as VQA-RAD lau2018dataset, PathVQA he2020pathvqa, and OmniMedVQA hu2024omnimedvqa, test text–image reasoning but often rely on templated questions, limiting their coverage of authentic diagnostic challenges. Recent meta-benchmarks (e.g., Rad-Bench kuo2024rad, DiagnosisArena zhu2025diagnosisarena, MedAgentsBench tang2025medagentsbench) have emphasized comprehensiveness, yet remain breadth-oriented: they focus on scale and diversity rather than probing reasoning under uncertainty.
@@ 2.2 VLM Evaluation Methodologies
Medical VLMs are typically evaluated via three strategies: reference-based, reference-free, or hybrid metrics. Traditional reference-based methods include BLEU, ROUGE, and CIDEr for surface overlap, as well as semantic similarity scores such as BERTScore zhang2019bertscore, CLIPScore hessel2021clipscore, and domain-specific adaptations like CheXbert smit2020chexbert or RadGraph-F1 jain2021radgraph. While scalable, these methods cannot assure that generated outputs are clinically grounded. Reference-free approaches, such as saliency attribution (e.g., Grad-CAM selvaraju2020grad) or image–text alignment tests, serve as indirect proxies for reasoning but lack explanatory power. More recent frameworks have introduced LLM-based grading systems for evaluating factual accuracy and reasoning plausibility in free-form clinical responses. For example, the Expert-of-Experts Verification and Alignment (EVAL) framework streamlines clinician-led evaluation in bleeding scenarios smith2025eval, while LLMEval-Med combines expert-designed checklists with LLM-as-Judge dynamics to improve reliability in medical QA and reasoning tasks zhang2025llmevalmed; zhi2025medgr. Despite their automation advantages, these approaches remain subject to ongoing reliability validation.
@@ 4.1 Models Benchmarked
We benchmark three categories of models to provide a comprehensive view of current VLM capabilities. First, we include proprietary frontier systems with strong multimodal reasoning ability, namely GPT-4o openai2024gpt4o and Gemini comanici2025gemini, which serve as reference baselines. Second, we evaluate powerful open-source generalist models such as Claude anthropic2024claude, Doubao doubao2024, Qwen-VL-Plus bai2023qwen. Finally, to assess the effect of domain-specific adaptation, we benchmark medical-specialized VLMs including LLaVA-Med Li2023LLaVAMed, RadFM Wu2023RadFM, Med-Flamingo moor2023medflamingo and MedGemma sellergren2025medgemma, which represent current efforts to tailor VLMs for clinical applications. A full list of models, versions, and hyperparameters is provided in Appendix C.
@@ 7 Ethics Details
  • Radiopaedia: Radiopaedia data used in this study are provided under the Creative Commons Attribution-NonCommercial-ShareAlike 3.0 (CC-BY-NC-SA 3.0) license. The data were utilized for non-commercial purposes with proper attribution to Radiopaedia and are redistributed under the same licensing terms.